OpenStudy (anonymous):

Astronomers estimate that a 2 - km wide asteroid collides with the Earth once every million years. The collision could pose a threat to life on Earth. (a) Assume a spherical asteroid has a mass of 3200 kg for each cubic meter of volume and moves toward the Earth at 15 km/s. How much destructive energy could be released when it embeds itself in the Earth? (b) For comparison, a nuclear bomb could release about . How many such bombs would have to explode simultaneously to release the destructive energy of the asteroid collision with Earth?

OpenStudy (anonymous):

so we are given that the astroid is spherical and has a width (aka diameter) of 2km wide. we will break the question up into steps. first, find the volume of the astroid, using the equaiton of a sphere \[\frac{ 4 }{ 3 }\times \pi \times r ^{3} = volume\] where r is the radius of the sphere